# Catalog-Native LLM: Speaking Item-ID dialect with Less Entanglement for Recommendation

### v1

@@ Abstract
While collaborative filtering delivers predictive accuracy and efficiency, and Large Language Models (LLMs) enable expressive and generalizable reasoning, modern recommendation systems must bring these strengths together. Growing user expectations, such as natural-language queries and transparent explanations, further highlight the need for a unified approach. However, doing so is nontrivial. Collaborative signals are often token-efficient but semantically opaque, while LLMs are semantically rich but struggle to model implicit user preferences when trained only on textual inputs. This paper introduces Item-ID + Oral-language Mixture-of-Experts Language Model (IDIOMoE), which treats item interaction histories as a native dialect within the language space, enabling collaborative signals to be understood in the same way as natural language. By splitting the Feed Forward Network of each block of a pretrained LLM into a separate text expert and an item expert with token-type gating, our method avoids destructive interference between text and catalog modalities. IDIOMoE demonstrates strong recommendation performance across both public and proprietary datasets, while preserving the text understanding of the pretrained model.
@@ 1 Introduction
Conventional recommenders like collaborative filtering (CF)(Koren et al., 2009), content-based (CB)(Lops et al., 2011), and sequential models (Kang & McAuley, 2018; Sun et al., 2019; Zhai et al., 2024) perform well within their scope when data are abundant, but they depend heavily on the quality of logs and item attributes. They remain vulnerable to popularity bias (Abdollahpouri et al., 2019), struggle to integrate heterogeneous signals (text, behavior, and context), and cannot support natural language queries.
@@
Inspired by mixture-of-experts (MoE) (Shazeer et al., 2017; Lepikhin et al., 2020; Fedus et al., 2022), we view ItemID modeling as a dialect distinct from natural language. But unlike standard MoE, which routes tokens indiscriminately, we design a targeted Item-ID + Oral-language Mixture-of-Experts Language Model (IDIOMoE) that assigns a dedicated collaborative expert for IDs alongside a preserved text expert for language. A token-type gate orchestrates their interaction, mitigating interference while retaining pretraining knowledge. Evaluations on both public benchmarks and a real-world industrial dataset from a leading online platform with hundreds of millions of users show that IDIOMoE consistently outperforms text-only adapters and item-only baselines. Our main contributions are:
@@ 2.1 Conventional Recommendation Methods
Traditional recommendation models fall into collaborative filtering (CF), content-based (CB), and sequential paradigms. CF learns from user–item interactions to model latent preferences (Koren et al., 2009), while CB leverages item attributes to improve personalization and mitigate cold-start issues (Lops et al., 2011). Sequential models further capture temporal dynamics, using models such as RNNs (Hidasi et al., 2015), SASRec (Kang & McAuley, 2018), and BERT4Rec (Sun et al., 2019). Though these models achieve strong performance under sufficient data, they operate on opaque ID sequences and require hand-crafted features or specialized architectures to incorporate diverse signals like language or intent. They also struggle with long-tail exposure (Abdollahpouri et al., 2019).
@@ 2.2 Generative Recommendation
Some works treat recommendation as sequence generation, unifying retrieval and ranking under a generative objective (Yang et al., 2025). This includes large-scale decoder models such as HSTU (Zhai et al., 2024), which scales to trillions of parameters, and OneRec (Deng et al., 2025), which uses a sparse MoE encoder–decoder architecture for scalable training. These approaches improve novelty, fluency, and explainability, but are resource-intensive and can forget collaborative structure if not carefully aligned with interaction data. They also do not support conversational recommendation.
@@ 2.3 Motivation and Positioning
We address this challenge by introducing a Item-ID + Oral-language Mixture-of-Experts Language Model (IDIOMoE) that treats item interactions as a native dialect. IDIOMoE dedicates separate pathways to item and text processing in each block, with a lightweight token-type gate that reduces interference while retaining language understanding. This design enables efficient ID-level modeling and better alignment with both semantic and collaborative objectives.
@@ 3.2 IDIOMoE
We present the Item-ID + Oral-language Mixture-of-Experts Language Model (IDIOMoE), a pretrained decoder-only LLM augmented with item-specialized experts and native item tokens. IDIOMoE keeps the language skills of the base model intact while learning collaborative patterns directly from user-item sequences. We start from a pretrained causal transformer and replace each feed-forward network (FFN) with a two-expert module:
@@
Routing is handled by a static token-type gate: tokens that are item IDs are sent to the item expert; all other tokens go to the text expert. Only one expert is active per token, so compute stays comparable to the base model (Appendix B.6.3). Attention and normalization layers are shared among all token types. Figure 2 provides an overview of our framework.
@@ 4.1 Experimental Settings
Our main focus is on LLM-based recommenders, so the most relevant baselines are different ways of adding recommendation capability to LLMs. We include established LLM-for-Rec baselines that are directly comparable to our setting: the P5/P5-CID family, which reframes recommendation as text-to-text generation over a pretrained language model (Geng et al., 2022; Hua et al., 2023); VIP5, a multimodal extension of P5 that adapts the LLM with parameter-efficient modules (Geng et al., 2023); E4SRec, which keeps the LLM largely frozen and adds a lightweight ID-side adapter for sequential recommendation (Li et al., 2023d); and ReAT, which aligns LLMs to recommendation objectives via auxiliary, recommendation-specific generated tasks (Cao et al., 2024). These capture the main design choices for adding recommendation capability to LLMs (prompting, adapters, frozen-backbone adapters, alignment), and thus form our most relevant comparison set. In addition, we compare three variants built on the same backbone: (i) ID Transformer, trained only on item tokens; (ii) Item-ID LLM + text-derived bias (Jiang et al., 2025), where ID embeddings are augmented with text features; and (iii) Item-LLM, which integrates item text via vocabulary expansion but without MoE. These three variants are matched to IDIOMoE in parameter count and trained under identical token budgets. For completeness, we also report results of classical sequence models (GRU4Rec (Hidasi et al., 2015), Bert4Rec (Sun et al., 2019), FDSA (Zhang et al., 2019), S3-Rec (Zhou et al., 2020)), recent quantized/contrastive approaches (VQ-Rec (Hou et al., 2023b), MissRec (Wang et al., 2023a), TIGER (Rajput et al., 2023), MQL4GRec (Zhai et al., 2025), IDGenRec (Tan et al., 2024)), and strong transformer baselines (SASRec (Kang & McAuley, 2018), HSTU (Zhai et al., 2024)). We further include CoVE (Zhang et al., 2025), which extends an LLM with LoRA parameters to encode catalog items. While these embedding-driven or classical models are not our primary comparison targets, we include them for completeness on smaller Amazon datasets. Full baseline details are in Appendix B.1.
@@ 4.1.1 Results: Amazon Catalogs
We evaluate SASRec (Kang & McAuley, 2018), HSTU (Zhai et al., 2024), ID-Transformer, LLM-based baselines and IDIOMoE on Larger Amazon datasets. Table 4.1.1 presents the results. IDIOMoE is the strongest LLM-based method across all three catalogs: it is the top LLM on Beauty (2nd overall behind HSTU by a small margin), and it achieves the best overall scores on Books and Toys. In contrast, Item-LLM and Text-Attr LLM Jiang et al. (2025) lag behind IDIOMoE across metrics, indicating that simply mixing item/text tokens or adding text-derived biases is insufficient. Together, these results support our claim that disentangling item and language pathways yields higher ranking quality than prior LLM baselines while remaining competitive with the best non-LLM models.
@@ 4.1.2 Results: Proprietary Dataset
Figure 4.1.1 (Table B.6.1) shows results on our large-scale proprietary dataset as improvements over the SASRec Kang & McAuley (2018) baseline. ID-Transformer achieves strong gains, confirming that transformers can effectively capture collaborative filtering signals when grounded in IDs and given enough compute. Title-LLM, which relies solely on free-form item titles, collapses in performance, highlighting the limitations of text-only representations for recommendation. Item-LLM combines IDs with textual features and yields further improvements, particularly on HR@10, demonstrating the value of jointly modeling collaborative and semantic signals. HSTU Zhai et al. (2024) provides modest gains but falls short compared to the LLM-based approaches and doesn’t support explainable recommendation. Finally, our method (IDIOMoE) achieves the largest improvements across all metrics (+27.1% NDCG@10, +16.6% HR@10, +31.2% MRR), showing that disentangling ID and text processing with specialized experts and routing not only preserves collaborative filtering strength but also better leverages semantic cues for robust large-scale recommendation.
@@ 4.2 Ablations
We also compare with various MoE designs. Both MoA (expert attention modules) and MoT (expertized full transformer blocks with cross attention) yield large improvements over all non-MoE controls. Importantly, IDIOMoE performs on par or better than both, despite using a simpler and more efficient expert design focused solely on FFNs with static routing.
@@ 4.2.3 Where to Insert MoE Layers
To study where MoE layers are most effective, we conduct an ablation by selecting different insertion strategies. Specifically, we activate MoE experts in (i) the first 8 layers, (ii) the middle 8 layers, (iii) the last 8 layers, and (iv) every third layer throughout the model. This allows us to compare the impact of placing MoE capacity in shallow, intermediate, deep, or evenly distributed positions. We report results on the Amazon-Arts dataset in Table 4.2.3. We observe clear differences depending on where MoE layers are inserted. Using MoE in the first 8 layers yields the weakest performance, suggesting that early representations are dominated by low-level token processing where additional capacity is less beneficial. Distributing MoE every three layers achieves moderate improvements but still falls short. Placing MoE in the middle 8 layers improves results, but the largest gains come from inserting MoE in the last 8 layers (+27.6% HR@10 and +28.4% NDCG@10 over baseline). This indicates that deeper layers (where task-specific semantics and collaborative filtering patterns are most prominent) benefit most from specialized experts, as they directly shape the final ranking representations.
@@ A.1 Classic Recommendation Approaches
Recommender systems have long relied on two complementary paradigms: collaborative filtering (CF) (Yao et al., 2021; Wang et al., 2025; Li et al., 2022; He & McAuley, 2015) and content-based (CB) methods. CF models exploit user–item interaction patterns, such as ratings or clicks, to learn latent representations of users and items (Koren et al., 2009). This approach is domain-agnostic and often yields high accuracy, but it suffers from well-known cold-start problems for new users or items and can exhibit strong popularity bias (Abdollahpouri et al., 2019), over-recommending popular items at the expense of long-tail discovery. CB methods instead leverage explicit item features or descriptions to recommend similar items, which can address item cold-start but ignore collaborative patterns and the “wisdom of the crowd.” These methods may produce over-specialized recommendations that limit serendipity.
@@
Moving beyond static recommendation, sequential models (Yuan et al., 2018; Zhou et al., 2020; de Souza Pereira Moreira et al., 2021; Hou et al., 2022; 2023a; Wang et al., 2023a) predict a user’s next interaction by modeling temporal dependencies in their history. Early neural solutions include GRU4Rec (Hidasi et al., 2015), which applied gated recurrent units to capture sequence dynamics. The introduction of Transformers brought a step-change: SASRec (Kang & McAuley, 2018) was the first to model next-item prediction in an autoregressive fashion using self-attention, improving short-term preference modeling. BERT4Rec (Sun et al., 2019) adapted bidirectional Transformers to better utilize context on both sides of a target position. These architectures form strong baselines in academic and industrial settings, yet they still rely on abstract IDs or dense embeddings, making it hard to integrate external semantic knowledge or to directly optimize multiple objectives beyond accuracy.
@@ A.2 Large Language Models for Recommendation
The advent of large language models (LLMs) pretrained (Yuan et al., 2020; Xiao et al., 2021; Qiu et al., 2021; Li et al., 2021; Yuan et al., 2021; Shin et al., 2022) on massive corpora has opened new opportunities for recommendation. (Zeng et al., 2020; Liu et al., 2023b; Lin et al., 2024; Yuan et al., 2023; Wang et al., 2024a; Fu et al., 2024) LLMs provide broad world knowledge, reasoning skills, and instruction-following Zhang et al. (2023); Li et al. (2024a); Contal & McGoldrick (2024) abilities that can extend beyond the pattern-matching of traditional recommenders (Zhang et al., 2021b; Muhamed et al., 2021; Cui et al., 2022; Liu et al., 2022; Zhang & Wang, 2023; Wei et al., 2024; Li et al., 2023b; Wang et al., 2023b).
@@
A pioneering example is P5 (Geng et al., 2022), which reformulates diverse recommendation tasks into a unified text-to-text format, allowing zero-shot Hou et al. (2024b) and few-shot transfer between tasks such as rating prediction, sequential recommendation, and explanation generation (Bao et al., 2023a; Li et al., 2023c; Yue et al., 2023; Lu et al., 2023; Zhang et al., 2021a; Wu et al., 2024). This unification facilitates integration of multiple modalities, such as textual descriptions or reviews, and enables natural-language queries Liu et al. (2023a); Bao et al. (2023b); Dai et al. (2023); Lin & Zhang (2023); Zhang & Wang (2023); Yang et al. (2023); Carranza et al. (2024); Kieu et al. (2025). However, item representation in such setups is often token-inefficient—especially for large catalogs—because items must be described in text, and off-the-shelf LLMs lack direct exposure to collaborative signals from user–item interactions (Cao et al., 2024). This leads to a mismatch between the LLM’s pretrained knowledge and the domain-specific collaborative knowledge needed for effective recommendation.
@@
To address these limitations, fine-tuning methods adapt LLMs to recommendation tasks while preserving language capabilities Ren & Huang (2024); Zhao et al. (2025); Li et al. (2024b); Wang et al. (2021). GDM (Cao et al., 2024) introduces auxiliary natural-language training tasks (e.g., masked item modeling, BPR) to inject collaborative patterns. MQL (Zhai et al., 2025) encodes multimodal item attributes (text, images) into a shared quantitative token space, enhancing cold-start and cross-domain performance. RL-based alignment (Lu et al., 2024) further improves controllability by optimizing instruction-following behavior with preference-based rewards, enabling conversational Friedman et al. (2023); Li et al. (2019); Chen et al. (2019); Kemper et al. (2024); Li et al. (2023a); Tang et al. (2025) and constraint-aware recommendation.

### v2

@@ Abstract
While collaborative filtering delivers predictive accuracy and efficiency, and Large Language Models (LLMs) enable expressive and generalizable reasoning, modern recommendation systems must bring these strengths together. Growing user expectations, such as natural-language queries and transparent explanations, further highlight the need for a unified approach. However, doing so is nontrivial. Collaborative signals are often token-efficient but semantically opaque, while LLMs are semantically rich but struggle to model implicit user preferences when trained only on textual inputs. This paper introduces Item-ID + Natural-language Mixture-of-Experts Language Model (IDIOMoE), which treats item interaction histories as a native dialect within the language space, enabling collaborative signals to be understood in the same way as natural language. By splitting the Feed Forward Network of each block of a pretrained LLM into a separate text expert and an item expert with token-type gating, our method avoids destructive interference between text and catalog modalities. IDIOMoE demonstrates strong recommendation performance across both public and proprietary datasets, while preserving the text understanding of the pretrained model.
@@ 1 Introduction
Conventional recommenders like collaborative filtering (CF)(Koren et al., 2009), content-based (CB)(Lops et al., 2011), and sequential models (Kang and McAuley, 2018; Sun et al., 2019; Zhai et al., 2024) perform well within their scope when data are abundant, but they depend heavily on the quality of logs and item attributes. They remain vulnerable to popularity bias (Abdollahpouri et al., 2019), struggle to integrate heterogeneous signals (text, behavior, and context), and cannot support natural language queries.
@@
Inspired by mixture-of-experts (MoE) (Shazeer et al., 2017; Lepikhin et al., 2020; Fedus et al., 2022), we view ItemID modeling as a dialect distinct from natural language. But unlike standard MoE, which routes tokens indiscriminately, we design a targeted Item-ID + Natural-language Mixture-of-Experts Language Model (IDIOMoE) that assigns a dedicated collaborative expert for IDs alongside a preserved text expert for language. A token-type gate orchestrates their interaction, mitigating interference while retaining pretraining knowledge. Evaluations on both public benchmarks and a real-world industrial dataset from a leading online platform with hundreds of millions of users show that IDIOMoE consistently outperforms text-only adapters and item-only baselines. Our main contributions are:
@@ 2.1 Conventional Recommendation Methods
Traditional recommendation models fall into collaborative filtering (CF), content-based (CB), and sequential paradigms. CF learns from user–item interactions to model latent preferences (Koren et al., 2009), while CB leverages item attributes to improve personalization and mitigate cold-start issues (Lops et al., 2011). Sequential models further capture temporal dynamics, using models such as RNNs (Hidasi et al., 2015), SASRec (Kang and McAuley, 2018), and BERT4Rec (Sun et al., 2019). Though these models achieve strong performance under sufficient data, they operate on opaque ID sequences and require hand-crafted features or specialized architectures to incorporate diverse signals like language or intent. They also struggle with long-tail exposure (Abdollahpouri et al., 2019).
@@ 2.2 Generative Recommendation
Some works treat recommendation as sequence generation, unifying retrieval and ranking under a generative objective (Yang et al., 2025). This includes large-scale decoder models such as HSTU (Zhai et al., 2024), which scales to trillions of parameters, and OneRec (Deng et al., 2025b), which uses a sparse MoE encoder–decoder architecture for scalable training. These approaches improve novelty, fluency, and explainability, but are resource-intensive and require careful objective and data design to fully exploit collaborative interaction signals. They also do not support conversational recommendation.
@@ 2.2.1 LLM-Based Recommendation and Semantic–ID Alignment
2.3 Multimodal MoE LLMs
Recent work integrates MoE into multimodal LLMs (MLLMs) and LVLMs Bao et al. (2022); Shen et al. (2023); Diao et al. (2025); Deng et al. (2025a). MoE-LLaVA (Lin et al., 2024a) adds a sparse MoE backbone to LLaVA (Liu et al., 2023a), converting feed-forward blocks into experts to match or exceed larger dense variants while activating fewer parameters and reducing visual hallucinations. Uni-MoE (Li et al., 2025) scales unified multimodal LLMs across many modalities and tasks with MoE layers. MoME (Xu et al., 2024) further mitigates task interference by factorizing the model into a Mixture of Vision Experts (MoVE) and a Mixture of Language Experts (MoLE), with MoVE aggregating multi-encoder vision features via an instruction-conditioned router and MoLE using sparsely gated adapter experts.
@@ 2.4 Motivation and Positioning
We address this challenge by introducing a Item-ID + Natural-language Mixture-of-Experts Language Model (IDIOMoE) that treats item interactions as a native dialect. IDIOMoE dedicates separate pathways to item and text processing in each block, with a lightweight token-type gate that reduces interference while retaining language understanding. This design enables efficient ID-level modeling and better alignment with both semantic and collaborative objectives.
@@ 3.2 IDIOMoE
We present the Item-ID + Natural-language Mixture-of-Experts Language Model (IDIOMoE), a pretrained decoder-only LLM augmented with item-specialized experts and native item tokens. IDIOMoE keeps the language skills of the base model intact while learning collaborative patterns directly from user-item sequences. We start from a pretrained causal transformer and replace each feed-forward network (FFN) with a two-expert module:
@@
Routing is handled by a static token-type gate: We use a simple static routing scheme: only item-ID tokens <|it-.|> are routed to the item expert, and all other tokens (titles, attributes, etc.) are routed to the text expert. All tokens share the same self-attention layers at every depth, so IDs and text always attend to each other, and the MoE split only affects the FFN sublayers, i.e., where ID- vs. text-specific information is stored. This design lets the model jointly reason over blended textual attributes and item IDs while allocating separate capacity for catalog structure. Moreover, since only one expert is active per token, so compute stays comparable to the base model (See Appendix B.6.4 for a discussion of efficiency results). Figure 2 provides an overview of our framework.
@@ 4.1 Experimental Settings
Our main focus is on LLM-based recommenders, so the most relevant baselines are different ways of adding recommendation capability to LLMs. We include established LLM-for-Rec baselines that are directly comparable to our setting: the P5/P5-CID family, which reframes recommendation as text-to-text generation over a pretrained language model (Geng et al., 2022; Hua et al., 2023); VIP5, a multimodal extension of P5 that adapts the LLM with parameter-efficient modules (Geng et al., 2023); E4SRec, which keeps the LLM largely frozen and adds a lightweight ID-side adapter for sequential recommendation (Li et al., 2023d); and ReAT, which aligns LLMs to recommendation objectives via auxiliary, recommendation-specific generated tasks (Cao et al., 2024). These capture the main design choices for adding recommendation capability to LLMs (prompting, adapters, frozen-backbone adapters, alignment), and thus form our most relevant comparison set. In addition, we compare three variants built on the same backbone: (i) ID Transformer, trained only on item tokens; (ii) Item-ID LLM + text-derived bias (Jiang et al., 2025), where ID embeddings are augmented with text features; and (iii) Item-LLM, which integrates item text via vocabulary expansion but without MoE. These three variants are matched to IDIOMoE in parameter count and trained under identical token budgets. For completeness, we also report results of classical sequence models (GRU4Rec (Hidasi et al., 2015), Bert4Rec (Sun et al., 2019), FDSA (Zhang et al., 2019), S3-Rec (Zhou et al., 2020)), recent quantized/contrastive approaches (VQ-Rec (Hou et al., 2023b), MissRec (Wang et al., 2023a), TIGER (Rajput et al., 2023), MQL4GRec (Zhai et al., 2025), IDGenRec (Tan et al., 2024)), and strong transformer baselines (SASRec (Kang and McAuley, 2018), HSTU (Zhai et al., 2024)). We further include CoVE (Zhang et al., 2025), which extends an LLM with LoRA parameters to encode catalog items. While these embedding-driven or classical models are not our primary comparison targets, we include them for completeness on smaller Amazon datasets. Full baseline details are in Appendix B.1.
@@ 4.1.1 Results: Amazon Catalogs
We evaluate SASRec (Kang and McAuley, 2018), HSTU (Zhai et al., 2024), ID-Transformer, LLM-based baselines and IDIOMoE on Larger Amazon datasets. Table 4.1.1 presents the results. IDIOMoE is the strongest LLM-based method across all three catalogs: it is the top LLM on Beauty (2nd overall behind HSTU by a small margin), and it achieves the best overall scores on Books and Toys. In contrast, Item-LLM and Text-Attr LLM Jiang et al. (2025) lag behind IDIOMoE across metrics, indicating that simply mixing item/text tokens or adding text-derived biases is insufficient. Together, these results support our claim that disentangling item and language pathways yields higher ranking quality than prior LLM baselines while remaining competitive with the best non-LLM models.
@@ 4.1.2 Results: Proprietary Dataset
Figure 4.1.1 (Table B.6.1) shows results on our large-scale proprietary dataset as improvements over the SASRec Kang and McAuley (2018) baseline. ID-Transformer achieves strong gains, confirming that transformers can effectively capture collaborative filtering signals when grounded in IDs and given enough compute. Title-LLM, which relies solely on free-form item titles, collapses in performance, highlighting the limitations of text-only representations for recommendation. Item-LLM combines IDs with textual features and yields further improvements, particularly on HR@10, demonstrating the value of jointly modeling collaborative and semantic signals. HSTU Zhai et al. (2024) provides modest gains but falls short compared to the LLM-based approaches and doesn’t support explainable recommendation. Finally, our method (IDIOMoE) achieves the largest improvements across all metrics (+27.1% NDCG@10, +16.6% HR@10, +31.2% MRR), showing that disentangling ID and text processing with specialized experts and routing not only preserves collaborative filtering strength but also better leverages semantic cues for robust large-scale recommendation.
@@ 4.2 Ablations
We also compare with various MoE designs. Both MoA (expert attention modules) and MoT (expertized full transformer blocks with cross attention) yield large improvements over all non-MoE controls. Importantly, IDIOMoE performs on par or better than both, despite using a simpler and more efficient expert design focused solely on FFNs with static routing. Although MoA and MoT are competitive on Amazon-Beauty and occasionally match or slightly exceed IDIOMoE there, we emphasize the industrial-scale results as our primary evidence. On this large setting, the FFN-based MoE of IDIOMoE consistently outperforms MoA/MoT variants. Nonetheless, the pattern we observe might be dataset-dependent. The core idea of IDIOMoE is to treat catalog items as first-class citizens and to separate where information about IDs and text is stored. All three MoE variants we ablate are consistent with this idea. Our choice to place MoE in the FFNs is guided by the stronger and more stable gains we see on the large-scale industrial dataset.
@@ 4.2.3 Where to Insert MoE Layers
To study where MoE layers are most effective, we conduct an ablation by selecting different insertion strategies. Specifically, we activate MoE experts in (i) the first 8 layers, (ii) the middle 8 layers, (iii) the last 8 layers, and (iv) every third layer throughout the model. This allows us to compare the impact of placing MoE capacity in shallow, intermediate, deep, or evenly distributed positions. We report results on the Amazon-Arts dataset in Table 4.2.3.
@@
We observe clear differences depending on where MoE layers are inserted. Using MoE in the first 8 layers yields the weakest performance, suggesting that early representations are dominated by low-level token processing where additional capacity is less beneficial. Distributing MoE every three layers achieves moderate improvements but still falls short. Placing MoE in the middle 8 layers improves results, but the largest gains come from inserting MoE in the last 8 layers (+27.6% HR@10 and +28.4% NDCG@10 over baseline). This indicates that deeper layers (where task-specific semantics and collaborative filtering patterns are most prominent) benefit most from specialized experts, as they directly shape the final ranking representations.
@@ 5 Conclusion
6 Acknowledgments
The authors would like to thank Zahra Miri for her assistance in preparing the figures.
This work was partially supported by NSF IIS 2347592, 2348169, DBI 2405416, CCF 2348306, CNS 2347617, RISE 2536663.
@@ A.1 Classic Recommendation Approaches
Recommender systems have long relied on two complementary paradigms: collaborative filtering (CF) (Yao et al., 2021; Wang et al., 2025a; Li et al., 2022; He and McAuley, 2015) and content-based (CB) methods. CF models exploit user–item interaction patterns, such as ratings or clicks, to learn latent representations of users and items (Koren et al., 2009). This approach is domain-agnostic and often yields high accuracy, but it suffers from well-known cold-start problems for new users or items and can exhibit strong popularity bias (Abdollahpouri et al., 2019), over-recommending popular items at the expense of long-tail discovery. CB methods instead leverage explicit item features or descriptions to recommend similar items, which can address item cold-start but ignore collaborative patterns and the “wisdom of the crowd.” These methods may produce over-specialized recommendations that limit serendipity.
@@
Moving beyond static recommendation, sequential models (Yuan et al., 2018; Zhou et al., 2020; de Souza Pereira Moreira et al., 2021; Hou et al., 2022; 2023a; Wang et al., 2023a) predict a user’s next interaction by modeling temporal dependencies in their history. Early neural solutions include GRU4Rec (Hidasi et al., 2015), which applied gated recurrent units to capture sequence dynamics. The introduction of Transformers brought a step-change: SASRec (Kang and McAuley, 2018) was the first to model next-item prediction in an autoregressive fashion using self-attention, improving short-term preference modeling. BERT4Rec (Sun et al., 2019) adapted bidirectional Transformers to better utilize context on both sides of a target position. These architectures form strong baselines in academic and industrial settings, yet they still rely on abstract IDs or dense embeddings, making it hard to integrate external semantic knowledge or to directly optimize multiple objectives beyond accuracy.
@@ A.2 Large Language Models for Recommendation
The advent of large language models (LLMs) pretrained (Yuan et al., 2020; Xiao et al., 2021; Qiu et al., 2021; Li et al., 2021; Yuan et al., 2021; Shin et al., 2022; Shirkavand et al., 2025a) on massive corpora has opened new opportunities for recommendation. (Zeng et al., 2020; Liu et al., 2023c; Lin et al., 2024b; Yuan et al., 2023; Wang et al., 2024a; Fu et al., 2024) LLMs provide broad world knowledge, reasoning skills, and instruction-following Zhang et al. (2023); Li et al. (2024a); Contal and McGoldrick (2024) abilities that can extend beyond the pattern-matching of traditional recommenders (Zhang et al., 2021b; Muhamed et al., 2021; Cui et al., 2022; Liu et al., 2022; Zhang and Wang, 2023; Wei et al., 2024; Li et al., 2023b; Wang et al., 2023b).
@@
A pioneering example is P5 (Geng et al., 2022), which reformulates diverse recommendation tasks into a unified text-to-text format, allowing zero-shot Hou et al. (2024b) and few-shot transfer between tasks such as rating prediction, sequential recommendation, and explanation generation (Bao et al., 2023a; Li et al., 2023c; Yue et al., 2023; Lu et al., 2023; Zhang et al., 2021a; Wu et al., 2024). This unification facilitates integration of multiple modalities, such as textual descriptions or reviews, and enables natural-language queries Liu et al. (2023b); Bao et al. (2023b); Dai et al. (2023); Lin and Zhang (2023); Zhang and Wang (2023); Yang et al. (2023); Carranza et al. (2024); Kieu et al. (2025). However, item representation in such setups is often token-inefficient—especially for large catalogs—because items must be described in text, and off-the-shelf LLMs lack direct exposure to collaborative signals from user–item interactions (Cao et al., 2024). This leads to a mismatch between the LLM’s pretrained knowledge and the domain-specific collaborative knowledge needed for effective recommendation.
@@
To address these limitations, fine-tuning methods adapt LLMs to recommendation tasks while preserving language capabilities Ren and Huang (2024); Zhao et al. (2025); Li et al. (2024b); Wang et al. (2021); Shirkavand et al. (2025c). GDM (Cao et al., 2024) introduces auxiliary natural-language training tasks (e.g., masked item modeling, BPR) to inject collaborative patterns. MQL (Zhai et al., 2025) encodes multimodal item attributes (text, images) into a shared quantitative token space, enhancing cold-start and cross-domain performance. RL-based alignment (Lu et al., 2024) further improves controllability by optimizing instruction-following behavior with preference-based rewards, enabling conversational Friedman et al. (2023); Li et al. (2019); Chen et al. (2019); Kemper et al. (2024); Li et al. (2023a); Tang et al. (2025) and constraint-aware recommendation.